Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS October 2011  Not logged in ELOG logo
Message ID: 77     Entry time: Wed Feb 22 13:37:11 2012
Author: Christian Irmler 
Project: BELLE 
Subject: Origami Module #4 
Run Number: run103 
Events: 100k 
StartTime: 22.02.2012 13:35:08 
EndTime: 22.02.2012 13:54:08  
Data: Crap 

Sr90, large scintillator,
HVret @ Origami closed
HVret @ SMU connected
Origami Mod. #4
HV=+-40V, Ibias=1,6µA
CO2 cooling on, Tin=-20.5°C, Tout=-21.5°C

I²C settings for cooled APV chips --> same as used in run010

e:\cvi-projects\apvdaq_fadc\origami6_V2_multi6_cooled.cfg

Lab measurement with new dark box but old, large scintillator

Wires between HV connectors ground and HVret conntact at the mambo were open --> no HVret conntection between module and SMUs!
Moreover the SMU Inv. Nr. 545-078 has either an unstable output or problems with the current measurement.

ELOG V3.1.4-966e3dd